OpenStudy (uri):

\[\int \frac{ln(1+x)}{1+x^2}dx\]

Parth (parthkohli):

Calculus... and Uri? =|

OpenStudy (anonymous):

there is no closed form in terms of elementary functions :(

OpenStudy (uri):

Yes :( but theres a nice substitution.

OpenStudy (uri):

Wait. I forgot the limits!

OpenStudy (anonymous):

ahh now we can do something with limits

OpenStudy (uri):

\[\int\limits_{0}^{1} \frac{\ln(1+x)}{1+x^2}dx\]

Parth (parthkohli):

Try using \(x = \tan(t)\)

OpenStudy (uri):

trig sub wont work :(

OpenStudy (uri):

its not that easy parth :)

OpenStudy (anonymous):

that will work..but, yes...not that easy

Parth (parthkohli):

@uri If you can copy questions from M.SE, then why can't I copy answers? ^_^

OpenStudy (uri):

It's a past exam question.

OpenStudy (uri):

The limits are infinity lol!

Parth (parthkohli):

HBA, how's life?

OpenStudy (uri):

I'm not my brother.

Parth (parthkohli):

I'm not my brother? :-|

Parth (parthkohli):

If I help you, would you stop asking other people to ask questions on your account?

Parth (parthkohli):

Answer in yes or no.

OpenStudy (uri):

This is my own question. WHy would someone else post a questoin on my acc? how obnoxious. Are you saying I can't ask a calculus question?

OpenStudy (anonymous):

ok guys :D, lets go with \(x=\tan \theta\) then we will have\[\int_{0}^{\pi/4} \ln(1+\tan \theta) \ d\theta\]\[\int_{0}^{\pi/4} \ln(\sin \theta+\cos \theta) \ d\theta-\int_{0}^{\pi/4} \ln(\cos \theta) \ d\theta\]

OpenStudy (shubhamsrg):

seems to be a mistake,denominaotr will be sec^2 (theta)

OpenStudy (anonymous):

that will be cancel out with dx=sec^2 dtheta :)

OpenStudy (shubhamsrg):

sorry :P

Parth (parthkohli):

@uri May I know where this question is from?

OpenStudy (anonymous):

now write \(\sin \theta+\cos \theta\) as\[\sin \theta+\cos \theta=\sqrt{2} \ \cos(\frac{\pi}{4}-\theta)\]

OpenStudy (uri):

Apparently this substitution would work \[x= \frac{1-u}{1+u}\] But I dont get how you would know it would work :/ @ParthKohli: This is a question from one of my old books.

terenzreignz (terenzreignz):

Look at all the attention... and interrogation @uri is getting :D What in the blazes did you do to rile such people, Uri? :)

Parth (parthkohli):

@uri Which book?

OpenStudy (uri):

Old notes that my teacher gave me. And people here, especially Parth thinks that I dont have the intellect to solve these integrals. So notnice.

OpenStudy (anonymous):

what am i doing here :D ?? \[\int_{0}^{\pi/4} \ln(\sqrt{2} \ \cos(\frac{\pi}{4}-\theta)) \ d\theta-\int_{0}^{\pi/4} \ln(\cos \theta) \ d\theta\]\[\]

OpenStudy (uri):

Do you think I'm in 11 grade without not knowing Math?

Parth (parthkohli):

@uri What is the name of your teacher?

terenzreignz (terenzreignz):

Okay, ambassadors... cool off ... and just focus on the question... which I have no idea how to begin :P

OpenStudy (uri):

I'm Jw is he is gonna do this on all the Questions i ask..

OpenStudy (uri):

if he is*

terenzreignz (terenzreignz):

You're Jw? Jw means...?

OpenStudy (anonymous):

and finally\[\int_{0}^{\pi/4} \ln(\sqrt{2}) \ d\theta+\int_{0}^{\pi/4} \ln( \cos(\frac{\pi}{4}-\theta)) \ d\theta-\int_{0}^{\pi/4} \ln(\cos \theta) \ d\theta=\frac{\pi}{8} \ln 2\]

OpenStudy (uri):

wow nice!

terenzreignz (terenzreignz):

got it :)

OpenStudy (uri):

which identity did you use to change the limit to pi/4?

hartnn (hartnn):

mukushla 's magic strikes again :D

OpenStudy (anonymous):

because\[\int_{0}^{\pi/4} \ln( \cos(\frac{\pi}{4}-\theta)) \ d\theta=\int_{0}^{\pi/4} \ln(\cos \theta) \ d\theta\]its a very nice point in integrating\[\int_{0}^{a} f(a-x) \ dx=\int_{0}^{a} f(x) \ dx\]thank u hartnn :)

OpenStudy (uri):

i got it. thank youu!

OpenStudy (anonymous):

welcome uri

OpenStudy (callisto):

I know I'm stupid. Can anyone tell me how to get \(\sin \theta+\cos \theta=\sqrt{2} \ \cos(\frac{\pi}{4}-\theta)\)?

OpenStudy (shubhamsrg):

@uri should explain you :)

OpenStudy (uri):

I'll give you a chance,Shubh.

OpenStudy (shubhamsrg):

really ? why? o.O

OpenStudy (uri):

Mai apne weak students ku kabhi bura feel nahi karate :P

terenzreignz (terenzreignz):

anybody mind translating that?^ in the time you guys took to insist that the other explain it, @Callisto would have gotten it by now XD

Parth (parthkohli):

@terenzreignz It's a dialogue from a movie

OpenStudy (shubhamsrg):

@uri I don;t know, explain please, I'll also get to leanr :)

OpenStudy (shubhamsrg):

learn*

OpenStudy (uri):

First learn grammar xD

OpenStudy (shubhamsrg):

what does " xD " mean ?

terenzreignz (terenzreignz):

This looks like a rift that dates back to way before I started on OpenStudy about 9 months ago... Guess I'll just sit back, drink my tea and watch you guys at it :D

OpenStudy (uri):

I'll explain you this,Shubh.

OpenStudy (shubhamsrg):

waiting.

OpenStudy (shubhamsrg):

okay, how many errors can one find here http://gyazo.com/3a581c345d853e199d6a93e5e473115c

OpenStudy (callisto):

\[\sqrt{2} [sin x (\frac{1}{ \sqrt2})+ cos x (\frac{1}{ \sqrt2})]\]\[= \sqrt{2} ( sin x sin y +cos x cos y)\]where y =45 = pi/4 So, \[sinx + cosx = \sqrt{2}cos(\frac{\pi}{4}-x)\]

hartnn (hartnn):

enough of your arguments. anything not related to the Q, should not be posted.

OpenStudy (shubhamsrg):

I just wanted her to explain that to @Callisto and me.

OpenStudy (uri):

Just use the substituion \[ x = \frac{1-u}{1+u}\]

OpenStudy (callisto):

\[x = \frac{1-u}{1+u}\] \[dx = \frac{-(1+u) - (1-u)}{(1+u)^{2}}du=-\frac{2}{(1+u)^{2}}du\]So, \[\int \frac{ln(1+x)}{1+x^2}dx = \int \frac{ln(1+\frac{1-u}{1+u})}{1+(\frac{1-u}{1+u})^{2}}(-\frac{2}{(1+u)^{2}}du)\]\[= -\int \frac{2ln(\frac{2}{1+u})}{(1+u)^2+(1-u)^2}du\]\[= -\int \frac{ln(\frac{2}{1+u})}{1+u^2}du\]\[= \int \frac{ln(1+u)-ln2}{1+u^2}du\]Hmm.. Am I on the right track?

OpenStudy (anonymous):

Nice approach :)

OpenStudy (callisto):

@uri suggested this approach, so I'm trying this...

OpenStudy (uri):

nope :( x = (1-u)/(1+u) = 2/(1+u)-1

OpenStudy (uri):

from that it will be easy.

OpenStudy (callisto):

If I use x = (1-u)/(1+u) = 2/(1+u)-1, I'm getting the same...

OpenStudy (uri):

yeah sorry! now add it

OpenStudy (uri):

from the last product.

OpenStudy (callisto):

Add it?

OpenStudy (uri):

ln(2) - ln(1+u)/(1+u^2) /(1+u^2) => 2I = ln(2)/(1+u^2) => ln(2)*pi/4

OpenStudy (uri):

imagine the limits are there lol

OpenStudy (callisto):

Hmmm...... x=1, u = 0, x=0, u = 1 \[\int_0^1 \frac{ln(1+x)}{1+x^2}dx\] \[= \int_1^0 \frac{ln(1+u)}{1+u^2}-\frac{ln2}{1+u^2}du\]\[= \int_1^0 \frac{ln(1+u)}{1+u^2}du-ln2(tan^{-1}u)|_1^0\]\[= -\int_0^1 \frac{ln(1+u)}{1+u^2}du-ln2(tan^{-1}u)|_1^0\] Just want to make sure, in this case, can we consider the variable u as a dummy variable?

OpenStudy (uri):

hm i dont know what youre doing there :/ but you were on the right track..i just posted the last line of the solution.

OpenStudy (uri):

forget about this integral; its a waste of time!

OpenStudy (callisto):

How do you get 2I = ... ?

OpenStudy (uri):

you add it..add the two integrals together.

OpenStudy (callisto):

But they are of different variables.

OpenStudy (uri):

have you done integration by reduction formulae?

OpenStudy (callisto):

No. Just split the fraction into two and integrate respectively.

OpenStudy (callisto):

If we can treat 'u' as a dummy variable, then I can get the answer, but the question is if it is correct to treat it as a dummy variable.

Parth (parthkohli):

The best thing here is that Callisto took the question seriously and wanted to learn.

OpenStudy (anonymous):

oh man... 2 months ago :)) :D

Parth (parthkohli):

Good times, good times.

OpenStudy (anonymous):

parth let me show you something interesting :) :D

OpenStudy (callisto):

Not fair, not fair.... mukushla only show Parth something interesting, but don't show us :(

OpenStudy (anonymous):

:) i wanted to show my first question on OS...i cant load it, low speed :D

OpenStudy (anonymous):

ahh here it is :) :D http://openstudy.com/users/mukushla#/updates/4fe355efe4b06e92b872179c

Parth (parthkohli):

Haha, yes! I remember that question too, honestly

Parth (parthkohli):

I didn't know I was guiding an epic user :-(

OpenStudy (anonymous):

:-D :-p

terenzreignz (terenzreignz):

curious now